OpenStudy (anonymous):

evaluate the integral fro 0 to 2 x/(1+x^2)dx

OpenStudy (anonymous):

make the substitution \[u=1+x^2\] giving \[du=2xdx\]

OpenStudy (anonymous):

might as well change the limits of integration while we are at is so we don't have to change back.

OpenStudy (anonymous):

\[u(0)=1\] \[u(2)=5\] now the integral is \[\int_1^5\frac{1}{u}du\]

OpenStudy (anonymous):

sorry i forgot to divide by 2

OpenStudy (anonymous):

\[u=1+x^2\] \[du=2xdx\] \[du=\frac{dx}{2}\] integral is \[\frac{1}{2}\int_1^5\frac{1}{u}du\]

OpenStudy (anonymous):

anti-derivative of \[\frac{1}{u}=ln(u)\] get \[\frac{1}{2}(\ln(5)-\ln(1))=\frac{1}{2}\ln(5)\]

OpenStudy (anonymous):

why u use ln(5)-ln(1).. i dont know how to work with that